On Jan. 21 we surveyed LMIRCam Pupil. Image of Pupil mirrors was too low and to the left. We traded the 0.375" half rounds on both sides of the NIC mount for 0.5625" half rounds on the SX side, leaving the 0.375" half rounds on the DX side. The center of the roof mirror now appeared to be at the center of the LMIRCam cold stop receptacle. We tilted the back side (toward the elevator) of NIC up slightly to bring the pupil images up to about the center of the LMIRCam cold stop. Final refinement was carried out by tip and tilting both SX and DX roof mirrors. Note: The May and Nov. 2011 alignment probably had NIC tilted too far to the DX side, and compensating with the DX roof mirror. By surveying the roof mirror this time, we believe we have an alignment which is more centered in the beamcombiner optics.

    Alignment Update, Sat night:

    We aligned to the cold stop nominal position. We should be aligning to the cold stop horizontal position. This requires moving the pupils up by another ~20% of their diameter.

    Recipe for Alignment:

    • Find position of large apertures for LMIRCam when they are horizontal.
    • Tilt NIC or roof mirrors to center images of pupil mirrors in cold stop.
    • Image secondary by using M band filter, t=0.15s. Take background while rotating tertiaries.
    • Center secondary image in pupil mirror by tilting ellipse mirrors.

    BCU power issues:

    controlled using java web interface at:

    All channels (D00-D05) should be "OFF" (red). This is due to a wiring mistake for the ADAM.

    Sun. Jan. 22

    Aligned NAC source. Wire A can be seen at row 512, column ~1000 when internal source is inserted. Slightly out of focus. NAC trans. stage is at positive limit.

    A: 2.8 Ohms 1.5 V, 0.55 A row 512 when at positive limit

    B: 6 Ohms 1.5 V, 0.25 A row 481 when at positive limit.

    -300,000 steps moves B from row 481 to row 512.

    On Sky:

    DX: IE 0, CA -60

    SX: IE 45, CA -5

    DX Bayside: -170, -212, -130

    Dx Camera Lens -35, 70

    SPC: A=4.82, B=3.78 mrad

    The nominal LMIRCam focus was 40 steps different than before.

    Recipe for star acquisition on fitst night:

    • Use a bright star at high air mass. Send preset.
    • su lbto
    • Use ptspiral -s 1 -p 15 -d 1 . This will survey a region of 140". Could also do -p 7.
    • Once star is found, check position with ptlist
    • ptmodify 1 IE 0 for DX. ptmodify 1 CA -60
    • ??
    Dual Aperture testing:
    When both sides are auhtorized, telescope attempts to arbitrate between optics and mount to only move the side requested:
    Example: ptmodify 1 IE 10 moves the right side while keeping left fixed.

    Found fringes. Setpoint now at 43950. Best focus at z=-92 (interesting proxy for pathlength!)

    Copointing achieved for
    left: IE=-30 CA=+45. right: IE =0 CA=-55

    Took NRM data. Setpoint moved to 44,400 after moving to new target.

    Remember to use Firefox for SPC and translation motion.

    Mon. Jan 23

    Closed for storm.
    Testing NOMIC array. Top of array appears light sensitive but not row clocking. Bottom is n scale but not light sensitive.
    Checked voltages and currents. Noticed Vrow_lo is 1 V but test report says it was set at 0 V. Vdetcom was set to 1.5 V.

    NOMIC works!!. Vrow_lo needs to be at 0 v rather than 1.

    NOMIC is rotated 90 degrees CCW relative to LMIRCam detector.

    See misalignment between fixed stop and pupil wheel of ~60% of the diameter, perpendicular to the baseline.
    Measured defocus of pinhole, which is now focused on LMIRCam.
    Separation of spots is 29 pixels for pinhole. Separation is 21 pixels for pinhole grid. Implies focus is on window side of pinhole grid, while LMIRCam focus is on the other side of the pinhole grid.

    Tues. Jan. 24

    Checked which NOMIC stop was misaligned by putting paper in front of SPC mirror. Both masks appear to be correctly aligned in the direction of the baseline. Perpendicular to the baseline it is the fixed stop which is approximately aligned. This suggest the NOMIC wheel needs to be shifted by ~1 mm to the right on the image.

    Checked NIL alignment. Can't see clear image of pupil or image. Likely significant misalignment.

    BD positive=open, 1.5 M range
    Combined Nulling Mirror Slide: +=out of beam 2 M range.
    ID: += out of beam. 3 M range
    right shutter += open (?)
    left shutter += open


    /groups/lbti/search/index.rss?tag=hotlist/groups/lbti/search/?tag=hotWhat’s HotHotListHot!?tag=hot2/groups/lbti/sidebar/HotListguestGuest2015-07-06 19:14:01+00:002015-07-06 19:14:01updated3Added tag - hotguestGuest2015-07-06 19:09:13+00:002015-07-06 19:09:13addTag2First createdguestGuest2015-07-06 19:09:00+00:002015-07-06 19:09:00created1wiki2015-07-06T19:14:01+00:00groups/lbti/wiki/6bdffFalseExport Control Training/groups/lbti/wiki/6bdff/Export_Control_Training.htmlGuest3 updatesExport Control Training To take the UA Export Control training please go to: UA Export Control Training and under Required Training Courses select the link below ...Falseguest2015-07-06T19:14:01+00:00phinzPhil Hinz2014-10-05 06:41:02+00:002014-10-05 06:41:02updated7Added tag - hotphinzPhil Hinz2014-10-05 06:40:59+00:002014-10-05 06:40:59addTag6phinzPhil Hinz2014-10-04 03:51:08+00:002014-10-04 03:51:08updated5phinzPhil Hinz2014-10-04 03:48:32+00:002014-10-04 03:48:32updated4phinzPhil Hinz2014-10-04 03:43:21+00:002014-10-04 03:43:21updated3phinzPhil Hinz2014-10-04 03:28:04+00:002014-10-04 03:28:04updated2First createdphinzPhil Hinz2014-10-04 03:21:14+00:002014-10-04 03:21:14created1weblog2014-10-05T06:41:02+00:00groups/lbti/weblog/74d8cFalseLBTI October Run: Night 2/groups/lbti/weblog/74d8c/LBTI_October_Run__Night_2.htmlPhil Hinz7 updatesLBTI October Run: Night 2 Night 2 of our run was clear and calm. We continued Kepler followup in the first half of the night. Justin Crepp and his team were here to obs...Falsephinz2014-10-05T06:41:02+00:00hot/groups/lbti/search/index.rss?sort=modifiedDate&kind=all&sortDirection=reverse&excludePages=wiki/welcomelist/groups/lbti/search/?sort=modifiedDate&kind=all&sortDirection=reverse&excludePages=wiki/welcomeRecent ChangesRecentChangesListUpdates?sort=modifiedDate&kind=all&sortDirection=reverse&excludePages=wiki/welcome0/groups/lbti/sidebar/RecentChangesListmodifiedDateallRecent ChangesRecentChangesListUpdateswiki/welcomeNo recent changes.reverse5searchlist/groups/lbti/calendar/Upcoming EventsUpcomingEventsListEvents1Getting events…